OpenStudy (sphott51):

Jane plans to invest $500 at 8.25% interest, compounded continuously. After 14 years, how much money has she accumulated? Has her money doubled or tripled?

OpenStudy (ad7):

How much is it after the first year?

OpenStudy (sphott51):

I don't know? I have to use a formula

OpenStudy (ad7):

500+500 x 8.25 / 100 Which is also 500 (1+0.0825) The next year you get 500 (1+0.0825) (1+0.0825) = 500 (1+0.0825)^2

OpenStudy (ad7):

500 (1+0.085)^2=500 x 1.1718

OpenStudy (ad7):

After the 3rd year it is 500 (1+0.0825)^3, After 14 years it is 500 (1+0.085)^14

OpenStudy (sphott51):

so using this formula A= Pe^rt ?

OpenStudy (ad7):

Here in Austria we use different letters, but I think you got it